I need help with two factoring polynomial questions that should be simple.
1. x^2 - 5x
2. 9x^2 - 25


Answer by ikleyn(52864)

.
1.  x^2 - 5x = x*(x-5).


2. 9x^2 - 25 = (3x - 5)*(3x + 5).

Solved.